### v1

# CodeCloak: A Method for Evaluating and Mitigating Code Leakage by LLM Code Assistants

Abstract
LLM-based code assistants are becoming increasingly popular among developers. These tools help developers improve their coding efficiency and reduce errors by providing real-time suggestions based on the developer’s codebase. While beneficial, these tools might inadvertently expose the developer’s proprietary code to the code assistant service provider during the development process. In this work, we propose two complementary methods to mitigate the risk of code leakage when using LLM-based code assistants. The first is a technique for reconstructing a developer’s original codebase from code segments sent to the code assistant service (i.e., prompts) during the development process, enabling assessment and evaluation of the extent of code leakage to third parties (or adversaries). The second is CodeCloak, a novel deep reinforcement learning agent that manipulates the prompts before sending them to the code assistant service. CodeCloak aims to achieve the following two contradictory goals: (i) minimizing code leakage, while (ii) preserving relevant and useful suggestions for the developer. Our evaluation, employing GitHub Copilot, StarCoder, and CodeLlama LLM-based code assistants models, demonstrates the effectiveness of our CodeCloak approach on a diverse set of code repositories of varying sizes, as well as its transferability across different models. In addition, we generate a realistic simulated coding environment to thoroughly analyze code leakage risks and evaluate the effectiveness of our proposed mitigation techniques under practical development scenarios.
Index Terms:
Code assistant, LLM, DRL, Security, Privacy
§§footnotetext: Equal contribution
I Introduction
I-A Machine Learning-Based Code Assistant
The use of large language models (LLMs) for code generation is growing and becoming increasingly useful to the global community [34]. The efficiency of these models has been demonstrated in a number of studies [3]. Code assistance tools, like GitHub Copilot, [^1] https://github.com/features/copilot StarCoder, [^2] https://huggingface.co/bigcode/starcoder and Code Llama [^3] https://ai.meta.com/blog/code-llama-large-language-model-coding/ leverage advanced LLMs to provide real-time suggestions, improving developers’ coding efficiency and reducing errors.
The flow of interaction with a code assistant during the coding process takes place as follows:
  1. 1.
    The developer writes code or describes the desired functionality in code or free text, within an integrated development environment (IDE).
  2. 2.
    The code assistant client, which is often a plugin integrated in the IDE, actively monitors the code being developed. It gathers relevant contextual information, including the developer’s recent interactions in the IDE and segments of code that may be relevant from other code sections and files in order to generate code suggestions.
  3. 3.
    Then, based on the data gathered, the code assistant client generates a prompt that contains relevant information for generating useful suggestions and sends it to the code assistant service (also referred to as model).
  4. 4.
    The code assistant service uses an LLM to generate suggestions about how the code can be completed and sends them back to the code assistant client integrated in the IDE.
  5. 5.
    These suggestions are presented to the developer via the IDE interface. The developer can integrate suitable suggestions in the code.
I-B Privacy and Security Risks
While LLMs have become invaluable in various tasks, their integration and use may inadvertently introduce security risks and expose sensitive data. In the context of code assistants, the use of LLMs may introduce new vulnerabilities to the developed code [22] or unintentionally expose sensitive data (i.e., proprietary code).
Previous studies have addressed three main security and privacy concerns: (1) the risk of inferring whether specific code was used to train an LLM model, known as a membership inference attack [2, 33, 20]; (2) the challenge of preventing proprietary code from being recommended by an LLM-based code assistant, particularly when the model has been trained on similar code, such as code from an open-source repository, by employing code manipulation strategies before the code is released [11]; and (3) the exploration of whether an LLM code assistant can be poisoned during its training/fine-tuning phase, leading to the presentation of intentional suggestions of vulnerable code to developers, thereby introducing vulnerabilities into the developed software [23, 14, 10].
However, while these security and privacy concerns have received media attention, [^4] https://www.forbes.com/sites/siladityaray/2023/05/02/samsung-bans-chatgpt-and-other-chatbots-for-employees-after-sensitive-code-leak/?sh=581383d16078 to the best of our knowledge, the risk and potential extent of private code leakage via prompts transmitted to a third-party code assistant service, along with strategies to mitigate this risk, have not been investigated. We regard this as a significant risk, given that the entire codebase of a proprietary software of an organization could potentially be leaked to an untrusted third-party service provider or an adversary capable of intercepting communication between developers and code assistant services. Therefore, there is a need to thoroughly assess the potential extent of leakage via prompts and devise effective mitigation strategies.
One approach for addressing the privacy concern is to use a local code assistant LLM, which is possible due to the proliferation of open-source models. The use of local models prevents sending prompts outside the network. However, commercial remote alternatives, offer advantages in terms of model freshness, maintenance, performance, ease of use, and integration with well-designed plugins for IDEs. The convenience offered by these commercial services continues to attract developers, despite the privacy concerns associated with remote services.
Figure 1: An example of CodeCloak’s performance: On the left, we present an original prompt and the code suggested by StarCoder. On the right, we present the manipulated prompts resulted by applying CodeCloak and the code suggested by StarCoder for the manipulated prompt.
There is a lack of both comprehensive methods for evaluating the extent of such leakage and effective mitigation strategies that can minimize code exposure. In this research we address this gap and propose two complementary methods for mitigating the risk of code leakage.
First, we introduce a technique for reconstructing the developer’s original code from prompts sent from the the IDE to the code assistant service during the development process. This reconstruction method allows code leakage assessment and evaluation when using LLM-based code assistants.
Then, we present CodeCloak, a novel deep reinforcement learning (DRL) agent that manipulates code prompts before they are sent to the code assistant service. Such manipulations may include function removal, renaming variable or function names, or replacing code with textual summarization. CodeCloak aims to achieve two contradictory goals simultaneously: (i) minimizing code leakage and the exposure of the developer’s proprietary code (by introducing manipulations to the prompts), while (ii) preserving relevant and useful suggestions provided by the code assistant.
As an example, in Figure 1 we present an original prompt (left) and a manipulated prompt generated by CodeCloak (right), both provided as an input to the StarCoder code suggestion model. At the bottom, the code suggestion provided by StarCoder for the original prompt and manipulated prompt are presented (marked by the white box). As can be seen, while the manipulated prompt is significantly different from the original one, the response in both cases (suggested code) is similar.
We evaluated our method on various LLM-based code assistants, including CodeLlama and StarCoder, using a unique process simulating real coding environments. The results show that our reconstruction method achieves an average reconstruction success rate of over 90%, indicating the extent of the leakage and demonstrating the need to develop an effective code leakage mitigation mechanism. Our mitigation method, CodeCloak, successfully mitigates the leakage to an average of 44%, while maintaining an average similarity of 71% between the suggestions obtained for the manipulated prompts and the suggestions obtained for the original prompts.
I-C Contributions
The contributions of our work can be summarized as follows:
  • To the best of our knowledge, we are the first to explore the risk of code leakage via the prompts and propose a mitigation.
  • We present a method for reconstructing codebase from code segments; in our usecase, we utilize prompts sent from the IDE to the code assistant model. We utilize the proposed method in order to evaluate the amount and quality of the potential code leakage. However, this method is generic and can be utilized in other usecases, such as in forensics when analyzing code fragment artifacts.
  • We present CodeCloak, a novel method based on a DRL agent, which is aimed at both reducing code leakage and preserving useful suggestions. The DRL agent can be installed as a local agent. CodeCloak is generic and transferable and thus, can be applied to any LLM code assistant. This approach can also be extended to LLMs in other domains.
  • We provide to the research community a unique dataset containing authentic requests (prompts) and responses (suggestions) generated by simulating the code development process and interactions with code assistant service. This dataset captures the prompts sent to the code assistant services along with the corresponding suggestions received, enabling evaluation of code leakage and mitigation techniques.
II Deep Reinforcement Learning
Reinforcement learning (RL) is an area of machine learning (ML) that addresses multi-step decision-making processes. An RL algorithm normally consists of an agent that interacts with an environment in a sequence of actions and rewards. In each time step $t$, the agent selects an action $a_{i}\in A$ = $\{a_{1},a_{2},...,a_{k}\}$. As a result of $a_{i}$, the agent transitions from the current state $s_{t}$ to a new state $s_{t+1}$. The selection of the action may result in a reward, which can be either positive or negative. The goal of the agent in each state is to interact with the environment in a way that maximizes the sum of future rewards $G_{t}$. Eq. 1 represents the cumulative reward over time:
$G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$ (1)
where:
  • $G_{t}$ is the expected return at step $t$.
  • $R_{t+k+1}$ represents the reward received after $k$ steps from step $t$.
  • $\gamma\in[0,1]$ is the discount factor, which provides the present value to future rewards. A reward received $k$ steps in the future is valued at $\gamma^{k}$ times its immediate value.
In order to maximize Eq. 1, the agent needs to learn its optimal policy. The policy defines how the agent behaves at any given time; i.e., which action to select when in a given state. When faced with complex environments with high-dimensional data, it is common to use DRL, i.e., incorporate deep neural networks to learn the optimal policy. This enables the agent to learn more sophisticated strategies and make more informed decisions. DRL algorithms can efficiently explore vast state and action spaces and devise strategies for addressing complex problems. Moreover, DRL excels in handling partial and noisy information. Utilizing the reward function, DRL algorithms can reconcile multiple, sometimes conflicting objectives, making them invaluable in security-related scenarios.
III Related Work
In this section, we provide a review of related work relevant to our study, including work on: recent advancements in code reconstruction techniques, privacy and security risks inherent in LLM code assistants, and prompt editing.
III-A Code Reconstruction
Previous studies have primarily focused on reconstructing text documents or static codebases [29, 31]. However, research on dynamically reconstructing code segments through iterative development cycles, where code sections may become obsolete, overlap with other code segments, or undergo modifications, is lacking. Static approaches assume that all code segments are relevant; therefore, in the dynamic code reconstruction case, these approaches will fail and result in duplicated, and incorrect code. Our proposed code reconstruction method distinguishes itself by addressing those scenarios and aims to reconstruct the most relevant and up-to-date codebase from a given sequence of evolving code segments.
III-B Privacy and Security Risks
As the adoption of code assistant tools has increased, inherent security and privacy risks have emerged [28, 22]. Pearce et al. [22] investigated the security implications of GitHub Copilot’s code contributions by designing scenarios that crafted prompts sent to Copilot results in vulnerable code being suggested. Incorporating suggestions from code assistants can intentionally or unintentionally introduce vulnerabilities or malicious content into the developer’s software. This is a challenging case to detect, especially if the LLM code assistant has been compromised during their training or fine-tuning phase. Inan et al. [10] demonstrated how improperly use of these tools can perpetuate existing security flaws in their training data. Khoury et al. [14] evaluated 21 programs in various programming languages, focusing on their vulnerability to common security flaws. They found that the code often failed to meet minimal security standards.
There are also concerns regarding inference attacks, when code repositories are used to train LLMs. The risk of inferring whether specific code was used to train the LLM model, also know as a membership inference attack (MIA), was discussed in [2, 33, 20, 18]. For example, Lukas et al. [18] proposed a novel MIA methodology and evaluated it on multiple code generation LLMs. They designed specific prompts to induce personally identifiable information (PII) leakage in programming languages and established a practical, automated pipeline for verifying such leaks.
Ji et al. [11] addressed the challenge of protecting proprietary code against unauthorized learning by LLM-based code assistants, particularly when these models are trained on open-source repositories. They proposed a method that involves applying lightweight, semantics-preserving transformations to the code before it is open-sourced. This strategy effectively obstructs models trained on the modified code, safeguarding the code from unauthorized learning while preserving its readability and functionality.
Lin et al. [17] suggested a data level defense that converts sensitive information in the original LLM input prompt into an encrypted format that, while indecipherable to human interpretation, retains its informativeness for the LLM to process. This research focused on mitigating leakage for general LLMs by using emojis to replace text, while we focus on code LLMs.
Patil et al. [21] presented a model level defense, which deletes sensitive information from the LLM model weights directly, assuming access to the model parameters, while we assume black-box access to the LLM.
While the studies above focused on assessing the security and potential vulnerabilities of code suggested by the code assistant, or developing methods for protecting proprietary code, we focus on the risks of code leakage via prompts when using code assistance models.
III-C Prompt Editing
On the rise of pretrained generative language models, the domain of prompt engineering for these LLMs has rapidly gained traction. Consequently, there is a growing focus among researchers on prompt editing to refine and optimize input prompts to enhance model output quality and alignment with user intentions [35, 8]. Several studies have explored prompt editing techniques using RL. For example, the study performed by Zhang et al. [36] presented TEMPERA, an RL method for dynamically editing prompts to enhance LLM performance on NLP tasks (such as sentiment analysis and topic classification). Kong et al. [15] introduced PRewrite, an automated prompt engineering method using RL for prompt optimization. Our study introduces a data level defense that combines RL methods along with prompt editing, in order to mitigate the code leakage during the code writing (development) process. Whereas previous studies mainly focused on general LLMs in different phases of model deployment, our defense specifically aims at code assistant provider. Our methodology can be integrated into diverse LLM code assistant models, without access to the model itself or its parameters.
IV Code Reconstruction Method
Figure 2: An illustration of the proposed code reconstruction process long with its main components.
Figure 3: System and user prompts for code reconstruction using ChatGPT.
IV-A Problem Definition
In this section, we introduce our method for code reconstruction using prompts when using LLM-based code assistants. Our method aims to evaluate the extent of codebase leakage during the development process. In order to achieve this, three challenges must be addressed:
  1. 1.
    The need to determine an effective method for intercepting and monitoring prompts sent to the code assistant service during the development process;
  2. 2.
    The need to formulate and design a code reconstruction approach capable of reconstructing the developer’s codebase, based on LLMs, in which the prompts are aggregated over time; and
  3. 3.
    The need to define a method for evaluating the severity of the leakage, given the original and reconstructed codebases.
The main challenge in reconstructing code from the prompts sent to the code assistant over time stems from the frequent modifications made by the developer during coding, such as editing and deleting code segments in different locations. Because of such modifications, early prompts that have been leaked and captured may be irrelevant and outdated or duplicated, preventing the use of traditional reconstruction methods that combine all available code segments into one complete codebase [29, 31].
The main components of our proposed method are presented in Figure 2.
IV-B Data Collection and Preparation
To collect the prompts during the development process, the Data Leakage Monitor component (component 1 in Figure 2) intercepts the traffic between the IDE and the code assistant service and extracts the prompts that are being sent from the code assistant client to the service. This process is performed in real time, for example, by placing a proxy (such as Burp Suite [12]) between the IDE and the code assistant service.
In the next step, the proposed method performs a preparatory process, using the Data Preparation component (component 2 in Figure 2), which is focused on analyzing the prompts. Its goal is to extract the code segments and additional contextual data (for example, in some code assistants, the prompt includes the source files’ names from which the code segments were extracted). In cases where the prompt includes this information, we process the prompts by aggregating them by the file they belong to (based on the filename).Next, we order the prompts chronologically (by their sending timestamp). At this stage, for each file, we have an ordered sequence of code segments. Finally, we remove code segments that are contained within subsequent code segments to avoid duplication and reduce code reconstruction time. By removing redundant code segments, we streamline the reconstruction process, enabling out method to receive a more comprehensive and concise representation of the codebase segments, resulting in a more accurate code reconstruction.
IV-C Code Reconstruction
Following the data collection and preparation phase, our method employs an LLM, which serves as the Code Reconstruction component (component 3 in Figure 2). As input, this model receives the processed prompts in the order in which they were transmitted from the code assistant client (in the IDE) to the code assistant service. In case where the total segments’ size exceeds the max token limit the LLM is able to process, we send the LLM segments of each file separately. If the total segments’ size still exceeds the max token limit, we perform the following iterative process: we take the maximum possible segments (that do not exceed the max token limit) and send them to the LLM. In the next iteration, we provide the LLM the codebase reconstructed in the last step, along with new segments. We repeats this process until we complete processing all the segments.The Code Reconstruction component then combines these code segments, building a representation of the developer’s final code. When these prompts are submitted to the LLM (ChatGPT4 in our evaluation), they are accompanied detailed instructions that clearly define the task (see Figure 3). This is done to ensure that the model produces relevant and accurate output consisting of the name and content of each file identified, providing a comprehensive view of the leaked code.
IV-D Leakage Evaluation
After reconstructing the codebase from the collected code segments, our next step is to evaluate the effectiveness of the code reconstruction process and the amount of code leakage (component 4 in Figure 2). This involves assessing whether the proposed method accurately reconstructed the code in its original form and whether the reconstructed version clearly conveys the original intent and functionality of the code. This evaluation process allows our method to accurately determine the percentage of the developer’s code that was leaked. In our evaluation, we use the normalized edit distance metric, which measures how similar two texts are by calculating the least number of changes needed to transform one text into another, adjusted by their lengths. This offers a straightforward, scaled gauge for visually comparing text differences and clearly depicting variations or mistakes.
V CodeCloak: Code Leakage Mitigation Method
V-A Problem Definition
As mentioned in Section I, to receive relevant suggestions, code assistant services rely on developers to provide the code they are currently working on. Let $P$ be the sequence of prompts sent from the developer’s IDE to the code assistant service during the development process. Correspondingly, let $S$ be the sequence of suggestions received from the code assistant service for the prompts $P$. Our goal is to achieve two seemingly contradictory objectives simultaneously: (i) minimize the exposure and leakage of the developer’s proprietary code by changing and manipulating the prompts $P$ before sending them to the code assistant service, and (ii) preserve the relevance and usefulness of the suggestions the developer receives for the manipulated prompts, by ensuring that they are as similar as possible to $S$.
In the context of code assistants, we observe that the prompts provided to the code assistant often contain excessive information that may not be essential for generating meaningful and relevant code suggestions. This observation serves as the basis of our proposed method.
To leverage this insight, we introduce a trained DRL model, which is tasked with identifying information that is not essential or has a negligible effect on the suggestions’ quality for a given prompt, and manipulate the prompt accordingly. By selectively modifying and manipulating the prompt, our DRL model aims to minimize the exposure of sensitive or proprietary code while preserving the ability of the code assistant to provide useful and relevant suggestions.
V-B CodeCloak Description
We propose CodeCloak, a novel DRL agent that manipulates prompts $P$ before sending them to the code assistant service. As illustrated in Figure 5, CodeCloak is placed between the IDE and the code assistant service, intercepting and modifying the prompts to reduce code exposure while preserving the relevance of the assistant’s suggestions. It can run locally on the developer’s machine or in the organization’s gateway.
The prompt manipulation process is modeled as a Markov decision process (MDP), where the agent learns to select the optimal actions (the manipulations) in order to maximize a reward that aims to minimize code leakage and maintain suggestion relevance.
The key components of the DRL agent are described in the subsections below:
V-B1 States
To enable the agent to effectively choose the appropriate manipulation, we provide it with detailed information about the prompt it is currently processing. In each iteration, the agent determines which manipulation to apply by analyzing the state that captures this information. We incorporate the following key elements into the state representation:
Prompt Embedding. The primary element is the prompt that is being manipulated. Since prompts are textual data, we need to represent them in a way that captures their semantic meaning and context. To achieve this, we feed the prompt to a pretrained code encoder model and use the last hidden state as an embedding representation of the prompt. This embedding technique allows the agent to capture the prompt’s semantic information, which is essential for making informed decisions. In some cases, the prompt’s length might exceed the encoder’s maximum token limit. To address this, we split the prompt into segments. Each segment is passed through the encoder to obtain its embedding representation and is represented in a separate state. In other words, each prompt segment is manipulated individually.
Relative Location. In the state, we include a fragment that represents the segment’s relative location within the full prompt. This positional information is important for the agent ,enabling it to better understand the order and placement of the code segments within the full prompt.
Cursor Position. The prompt sent to the code assistant includes the developer’s cursor location, helping the code assistant understand where in the code the developer is currently focusing on. In the state, we also include the relative locations of the developer’s cursor position (divided into two fragments: one indicates the cursor’s relative line location, and the second indicates the relative location in the line). Including the cursor position in the state representation is important, since it provides contextual information about the part of the code the developer is currently working on and allows the agent to understand what areas of the code it should focus on to apply the manipulations (for example, changing code that is “close” to the cursor might affect the code assistant’s suggestions more than manipulations in other areas).
The agent processes these states sequentially, applying manipulations in a loop over the segments. This segmentation strategy allows the agent to navigate and refine extensive prompts effectively, ensuring comprehensive coverage and strategic manipulations.
More formally, let $p_{t}$ be a prompt after $t$ manipulations, which is divided into $n$ segments: $p_{t,0},...,p_{t,n-1}$. We calculate the state representation for the $p_{t}$ ($\sigma_{t}$) as follows:
$i=t\bmod n$ is the location of the current segment to be manipulated at time $t$. Given $p_{t,i}$, the segment in the $i$th location at time $t$, the state embedding is: $\sigma^{emb}_{p_{t,i}}=\mathcal{L}(p_{t,i})$, where $\mathcal{L}(\cdot)$ is the pretrained encoder producing the embedding. We represent the segment’s relative location as $\sigma^{loc}_{p_{t,i}}$ and the cursor relative position for $p_{t}$ as $\sigma^{cur}_{t}$.
The complete state representation $\sigma_{t}$ is the concatenation:
$\sigma_{t}=[\sigma^{emb}_{p_{t,i}},\sigma^{loc}_{p_{t,i}},\sigma^{cur}_{t}]$ (2)
Observation Normalization. Given that we used the final hidden states of pretrained LLM as observation representation, there might be slight differences across various samples. In order to deal with this, we keep a running calculation of the mean and standard deviation for these observations and normalize them prior to inputting them into the policy and value networks. Such normalization is a standard practice in RL, and we observed that it enhances the efficacy of our approach.
V-B2 Prompt Manipulation Process
Given an initial prompt $p$, the agent starts in the state representation $\sigma_{0}$ as described above. In each timestep $t$, the agent selects an action $a_{t}$ from the set of possible manipulations $\mathcal{A}$, such as deleting or inserting lines, renaming variables or functions, or modifying functions. After applying the action $a_{t}$ to the prompt $p_{t}$, the agent transitions to a new state $\sigma_{t+1}$ representing the manipulated prompt $p_{t+1}$.
Note that in each step, the agent manipulates one segment of the total prompt. However, a manipulation applied to one segment might influence other segments. For example, if a variable name is changed in one segment, the agent will replace that variable name in all other segments where that variable instance appears.
V-B3 Actions
We defined a set of manipulations (referred to as $\mathcal{A}$) designed to provide our agent with the flexibility to effectively manipulate the prompts. In each step, our agent can apply one of the actions listed in Table I.
TABLE I: CodeCloak’s code manipulation actions
Action Description
Detect and Replace PII Detect PII (e.g., names, API keys, emails) in the current code segment and replace them with random strings of a fixed length.
Change Random Lines Swap positions of two random lines present in the current code segment.
Delete random line Remove a random line from the current code segment.
Insert random line Insert a random line in the current code segment.
Delete Functions body Incrementally Incrementally remove function bodies from the current code segment and replace the removed body replaced with summarization comments (that were generated using a local summarization model).
Delete Function Bodies (Keep Last) Remove all function bodies except the last one in the current code segment and replace the removed bodies with summarization comments (that were generated using a local summarization model).
Delete Function Bodies Remove all function bodies present in the current code segment and replace them with summarization comments describing the original function bodies.
Delete Functions Incrementally Incrementally remove functions from the current code segment.
Change Function Names Replace all function identifiers in the current code segment with random strings of a fixed length (5 characters) consisting of letters and numbers.
Change variable names Replace all variable identifiers in the current code segment with random strings of a fixed length (5 characters) consisting of letters and numbers.
Change Argument Names Replace all argument identifiers in the current code segment with random strings of a fixed length (5 characters) consisting of letters and numbers.
Stop Manipulations Stop processing the current prompt and send it to the code assistant service without further manipulations.
V-B4 Rewards
The reward function $r_{t}$ is designed to capture the trade-off between minimizing code leakage and preserving suggestion relevance. During training, the agent interacts with the code assistant service, sending manipulated prompts and receiving suggestions. This allows computation of the reward $r_{t}$ based on the similarity between the original and manipulated suggestions, guiding the agent towards an optimal prompt manipulation policy. Note that the reward considers the similarity of a full prompt and the corresponding suggestion received for the prompt.
Since we aim to consider both visual and functional similarity, we use the edit distance and cosine similarity metrics to calculate the reward. For visual similarity, we use the normalized edit distance ($NED$). The edit distance ($ED$[11] is defined as the minimum number of operations required to transform one string into another. The $NED$ normalizes the edit distance to a range from zero to one by dividing it by the maximum length of the two strings. A higher $NED$ value indicates that the two code sections are not visually similar.
For the functionality similarity measurement, we use cosine similarity ($CS$) [4, 7, 25]. $CS$ measures how similar two vectors are by calculating the cosine of the angle between them in a multi-dimensional space. To measure the $CS$ between two code sections (prompts or suggestions), we convert the code sections to their embedding representation using an embedding function $\mathcal{L}$ and apply the $CS$ metric. However, as mentioned in Section V-B1, the prompt length might exceed the maximum token limit that the encoder can process. Therefore, in cases in which the a code section exceeds the limit, we first divide the code section into $n$ segments and calculate the embedding representation of each segments separately, using $\mathcal{L}$. Then, we concatenate all the embedded segments and apply PCA [19]. Then, we use the result as the input for the $CS$ metric.
Since the $CS$ measures similarity but the $NED$ measures distance, we define the normalized-distance CS ($NDCS$) to simplify the calculation:
$\small DNCS(c,c^{\prime})=\frac{1-CS(\mathcal{L}(c),\mathcal{L}(c^{\prime}))}{2}$ (3)
The $DNCS$ normalizes the $CS$ to a range from zero to one, with a higher value implying that the two source codes do not have similar functionality.
Let $p_{t}$ and $p_{t}^{\prime}$ denote the original and manipulated prompts at timestep $t$, respectively. Let $s_{t}$ and $s_{t}^{\prime}$ denote the suggestions generated by the code assistant model for $p_{t}$ and $p_{t}^{\prime}$. We define the visual difference score $VD$ as:
$\small VD(p_{t},p_{t}^{\prime},s_{t},s_{t}^{\prime})=\lambda_{1}\cdot NED(p_{t},p_{t}^{\prime})-\lambda_{2}\cdot NED(s_{t},s_{t}^{\prime})$ (4)
We defined the functional difference score $FD$ as:
$\small FD(p_{t},p_{t}^{\prime},s_{t},s_{t}^{\prime})=\lambda_{1}\cdot DNCS(p_{t},p_{t}^{\prime})-\lambda_{2}\cdot DNCS(s_{t},s_{t}^{\prime})$ (5)
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are hyperparameters for the trade-off between the prompts and the suggestions distance . We chose to use complementary values to balance the two components; therefore $\lambda_{2}=1-\lambda_{1}$
The total reward score is defined as:
$\small RScore(p_{t},p_{t}^{\prime},s_{t},s_{t}^{\prime})=\\ \gamma_{1}\cdot FD(p_{t},p_{t}^{\prime},s_{t},s_{t}^{\prime})+\gamma_{2}\cdot VD(p_{t},p_{t}^{\prime},s_{t},s_{t}^{\prime})$ (6)
where $\gamma_{1}>0$ and $\gamma_{2}>0$ are hyperparameters for each reward metric component. We chose to use complementary values to balance the two components; therefore $\gamma_{2}=1-\gamma_{1}$
Naturally, $RScore$ awards a positive reward when the difference between the original and manipulated prompts is greater than the difference between the original and manipulated suggestions. Conversely, it awards a negative reward when the difference between the prompts is smaller than the difference between the suggestions. The goal is to optimize $RScore$ for the final manipulated prompt.
In traditional RL, the objective is to optimize the accumulated reward over the entire MDP. However, in our prompt manipulation scenario, we are primarily interested in the performance of the final manipulated prompt. Therefore, instead of using the $RScore$ directly, we propose using the difference in the $RScore$ between successive manipulations as the immediate reward:
$R_{t}=\\ RScore(p_{t},p_{t^{\prime}},s_{t},s_{t^{\prime}})-RScore(p_{t-1},p_{t^{\prime}-1},s_{t-1},s_{t^{\prime}-1})$ (7)
Figure 4: Training the DRL agent.
Figure 5: Trained agent interactions.
Algorithm 1 CodeCloak Inference Stage
1:Input: original prompt $p$, CodeCloak policy $\pi_{\text{CodeCloak}}$
2:Output: CodeCloak’s suggestion $s^{\prime}$
3:$p_{0}\leftarrow p$
4:Divide prompt $p_{0}$ into $n$ segments $p_{0,0},\ldots,p_{0,n-1}$
5:$t\leftarrow 0$
6:while True do
7:    $i\leftarrow t\bmod n$
8:    $\sigma_{t}\leftarrow\text{state representation for }p_{t,i}$
9:    $a_{t}\leftarrow\pi_{\text{CodeCloak}}(\sigma_{t})$
[…]
11:         break
12:    else
13:         $p_{t+1,i}\leftarrow\text{apply $a_{t}$ on }p_{t,i}$
14:         $t\leftarrow t+1$
15:$s^{\prime}\leftarrow\text{CodeAssistant suggestion for }p_{t}=p_{t,0},\ldots,p_{t,n-1}$
16:$s^{\prime}\leftarrow\text{apply CodeTranslation on }s^{\prime}$
17:return $s^{\prime}$
Reward Normalization. Since our reward comprises two distinct metrics, the values generated for each exhibit variation across diverse ranges (even if normalized). In order to deal with this, we use the applying exponential moving average for each reward component. We observed that this enhances the efficacy of our approach.
V-C CodeCloak Training Phase
In order to produce an efficient agent, capable of manipulating the prompts in an optimal manner, a DRL algorithm is used to train the agent (see Figure 4).
The input to the DRL agent is a dataset of prompts from various projects, all generated by the code assistant client. During its training, the agent processes the prompts and iteratively selects an action to apply on the prompt until the ‘stop manipulation’ action is selected; eventually, this will enable the trained DRL agent to successfully manipulate unseen or complex prompts during the inference phase.
To assess the effectiveness of the DRL agent’s manipulations during the training process, the agent interacts with the code assistant model and receives the code suggestions by sending the manipulated prompts to the code assistant model. This allows the agent to measure the similarity between the suggestions received for the original and manipulated prompts, and calculate the Reward required by the learning algorithm to train the DRL agent. The output of this phase is a trained DRL agent capable of effectively manipulating code prompts sent to an LLM-based code assistant in a manner in which the privacy of the code owner and the effectiveness/relevance of the code assistant’s suggestions are preserved. Note that the training of the DRL agent is performed according to a policy; this policy defines the preferred balance between preserving privacy (preventing the leakage of sensitive code) and preserving the original code suggestions made by the code assistant.
V-D CodeCloak’s Inference Phase
Following the training phase, the Mitigating component, which includes the DRL agent, is placed between the IDE and the code assistant service (see Figure 5).
During coding, each prompt sent from the developer’s IDE to the code assistant service is processed by the DRL agent before it is sent. The agent manipulates the prompt, minimizing the leakage while preserving the suggestion’s relevance, and sends the manipulated prompts to the code assistant service.
In some cases, an additional component called the Code Translation component, which is part of the Mitigating module, is applied to the returning suggestions of the code assistant service before it is sent back to the developer’s IDE; for example, in cases in which the DRL agent changes the names of functions/variables in the prompts before sending them to the code assistant service, the Code Translation component maps the names of the functions/variables to the original ones so they will be aligned with the developer’s code.
The complete process of the CodeCloak agent for a given prompt in the inference stage is outlined in Algorithm 1.
In our research, we used recurrent proximal policy optimization (recurrent PPO) [30] as our RL model. We chose to use this model due to its effectiveness in handling sequential data and managing temporal dependencies, which align well with our innovative segmentation strategy for state representation. Recurrent PPO facilitates our objective to manipulate prompts efficiently while aiming to preserve the desired quality of the suggestions. Its capacity to maintain the context across segmented inputs enhances our model’s ability to learn and adapt, providing a solid foundation for our prompt manipulation tasks.
VI Evaluation
VI-A Developer Coding Simulation
We developed a Python-based simulator designed to mimic the realistic behavior of developers during the coding process. The simulator incorporated features such as introducing coding errors at random intervals, correcting those errors after writing additional lines of code, simulating pauses in typing, going back and forth in the file, and maintaining a developers’ average typing speed rate. This simulator was used to generate datasets for our evaluation of both our code reconstruction method and our code leakage mitigation method (CodeCloak). It operated within an IDE configured with code assistant plugins, allowing us to capture the prompts sent from the IDE to the code assistant servers and the corresponding suggestions received. The prompts and suggestions captured during these simulated coding sessions formed the datasets used to evaluate our reconstruction and mitigation methods across different code assistants, including StarCoder, GitHub Copilot, and CodeLlama models.
VI-B Evaluation Setup
Models. The following LLM-based code assistant models were used in our experiments: GitHub Copilot [5] and local versions of StarCoder [16] and CodeLlama-13B [26]. To evaluate our reconstruction method, we used the StarCoder and Copilot models. Since GitHub Copilot’s policy does not allow running automation processes that manipulate and investigate the model’s behavior, we could not use the CodeCloak agent to manipulate prompts sent to Copilot. Therefore, we trained CodeCloak using StarCoder and evaluated it using the StarCoder and CodeLlama models.
Datasets. For the reconstruction method, we evaluated our approach using a dataset constructed from 20 GitHub repositories. Each repository in our dataset contains 3-4 files, with the number of lines per file ranging from 30 to 900. To train and evaluate CodeCloak, we created a dataset by randomly sampling repositories in the Python language from the CodeSearchNet [9] dataset. For each repository, we simulated a coding process by navigating to random locations in the repository, copying a random number of code lines to the IDE, and simulating typing from this point for a short period of time (between 20-90 seconds), capturing the generated prompts. This process yielded a dataset of around 30K diverse prompts of various sizes, from around 500 repositories, allowing CodeCloak to learn how to manipulate prompts effectively. In assessing the effectiveness of CodeCloak, we used prompts containing more than four functions since they contains substantial amount of code that can be leaked, and allow the agent to apply manipulations.
Code Reconstruction Details. We employed BURP Suite [^5] https://portswigger.net/burp/documentation/desktop/tools/proxy as an intermediary proxy to capture code segments as our simulator inputted the code into the IDE. We executed the coding simulating process within the PyCharm IDE. To reconstruct a codebase from the code segments, we used ChatGPT 4.0 (1106-preview version).
CodeCloak Details. To train the DRL agent, we used the RecurrentPPO algorithm from the Stable-Baselines3 Contrib library [24]. The state embedding was obtained by using the last hidden state of the StarEncoder [^6] https://huggingface.co/bigcode/starencoder model. The maximum token limit for input prompts was set at 1,600 tokens. To accommodate prompts that exceed the maximum token limit of the StarEncoder model, we divided each prompt into two segments. Unless otherwise mentioned, the output response length was set to 20 tokens. In addition, unless otherwise specified, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ were both set to 0.5, giving equal importance to both of CodeCloak’s objectives. Finally, based on empirical evaluation, the hyperparameters $\gamma_{1}$ and $\gamma_{2}$ were set to 0.8 and 0.2, respectively.
VI-C Evaluation Metrics
To evaluate the performance of our methods, we used two metrics to measure the similarity between two code sections:
Normalized Edit Distance. As previously mentioned, the edit distance is a metric that computes the distance between two text sections, normalized by the length of the longer section. In our case, it was used to compare the differences between two code segments. The results are presented as a percentage between 0% and 100%.
ChatGPT. Recent studies have explored the use of ChatGPT for various code-related applications, highlighting its potential as a useful tool for code analysis. Methods using ChatGPT for automatic code documentation [13], and summarization and refinement [6, 32] tasks demonstrate its ability to understand code semantics. ChatGPT’s successful use in static code analysis and application security testing [1, 27], highlight its code “understanding” capabilities. Building on these prior works that highlight ChatGPT’s strengths in diverse code-related tasks, we employed ChatGPT as a metric for measuring the similarity between code segments, considering both functional and visual aspects. We utilized ChatGPT to evaluate CodeCloak effectiveness of manipulations using two metrics, where for each metric we created a dedicated prompt: LeakGPT - scores the degree of code leakage from 0% to 100%, given the original code segment and the version obtained by an adversary. A higher score implies greater leakage. SugSimGPT - Given the original prompt and two code suggestions (the original and the one received for the manipulated prompt), ChatGPT provides a similarity score between 0% and 100%. A higher score indicates greater similarity between the suggestions. For each prompt, we instructed ChatGPT to consider different factors such as functional similarity, visual similarity, and the intended purpose or code meaning.
VI-D Code Reconstruction Results
We evaluated our code reconstruction method on 20 GitHub repositories, measuring the extent of code leakage when developing with the GitHub Copilot and StarCoder code assistants. These repositories were divided into small (repositories with up to 550 lines) and large (repositories with more that 550 lines) repositories. The results presented in Table II demonstrate a remarkably high reconstruction rate with an overall average reconstruction rate of 90.55% for Copilot and 91.86% for StarCoder, when the original repositories are almost completely leaked. These results emphasize the importance of mitigation such as CodeCloak.
Interestingly, the leakage rates were higher for smaller repositories than they were larger ones; for small repositories the average reconstruction rate was 94.10%, and for large repositories the rate was slightly lower at 88.3% We believe this occurs, since for smaller repositories, the LLM has wider contextual knowledge, enabling it to better reconstruct the codebase from the segments.
TABLE II: Reconstruction results for 20 repositories (average)
Repository Size Code LLM Average Reconstruction Rate $\uparrow$
Small Repositories GitHub 95.20%
StarCoder 93.05%
Large Repositories GitHub 85.91%
StarCoder 90.72%
VI-E CodeCloak Results
Our Agent’s Effectiveness. We compare the results of our agent with a baseline model, where for each prompt, we randomly choose manipulations that correspond to the number of manipulations that CodeCloak applied on this prompt. The results demonstrate CodeCloak’s effectiveness in balancing code leakage mitigation and preserving suggestion relevance. We’ve noticed that our agent tends to employ more aggressive actions, such as ’Delete Functions Body Incrementally’, in the initial steps. This suggests that the agent prioritizes removing redundant code and reducing potential leakage by making significant modifications early in the process. Subsequently, the agent transitions to more fine-tuned adjustments, such as ’Change Variable Names’ and ’Change Function Names’ in the later steps.
For suggestion relevance, for the normalized edit distance metric CodeCloak maintained a lower value of 37.92% compared to 53.55% for the baseline, indicating better preservation of relevant suggestions. Additionally, CodeCloak scored 75.54% for SugSimGPT, much higher than the 48.5% achieved by the baseline, ensuring highly relevant suggestions.
Regarding code leakage, for the normalized edit distance metric CodeCloak achieved 61.1% prompt dissimilarity, higher than the 50.5% prompt dissimilarity obtained by the baseline, indicating a greater distance between the original and manipulated prompt and better code obfuscation. CodeCloak also attains 59.55% for LeakGPT, lower than the 65.4% obtained by the baseline, demonstrating reduced code leakage.
These results highlight CodeCloak’s ability to strike an effective balance between mitigating code leakage and preserving suggestion relevance.
TABLE III: Performance comparison of the random and CodeCloak approaches
Normalized Edit Distance ChatGPT Metrics
Prompts $\uparrow$ Suggestions $\downarrow$ LeakGPT $\downarrow$ SugSimGPT $\uparrow$
Random 50.5% 53.55% 65.40% 48.50%
CodeCloak 61.1% 37.92% 59.55% 75.54%
Prioritizing Suggestions. We trained another model that prioritizes the objective of preserving the suggestions relevance over mitigating the leakage. Specifically, we set $\lambda_{1}=0.3$ and $\lambda_{2}=0.7$. The results presented in Table V demonstrate how CodeCloak can be adjusted to different components weights, depending on the developer’s needs.
TABLE IV: Performance comparison of CodeCloak with different lambda parameters
Normalized Edit Distance ChatGPT Metrics
Prompts $\uparrow$ Suggestions $\downarrow$ LeakGPT $\downarrow$ SugSimGPT $\uparrow$
$\lambda_{1}=0.5,\lambda_{2}=0.5$ 61.1% 37.92% 59.55% 75.54%
$\lambda_{1}=0.3,\lambda_{2}=0.7$ 40.00% 25.15% 82.23% 85.43%
Max Tokens Parameter. Max tokens refers to the maximum suggestion’s length that the code assistant generates for a given prompt. When training CodeCloak using Max tokens value 20, we evaluate the model’s performance in two different setting: $Maxtokens=20$ and $Maxtokens=50$. As can bee seen in Table IV, our agent successfully handles both settings, demonstrating CodeCloak’s ability to apply insights it acquired during the training in diverse scenarios.
TABLE V: Performance comparison of CodeCloak with different max tokens
Normalized Edit Distance ChatGPT Metrics
Prompts $\uparrow$ Suggestions $\downarrow$ LeakGPT $\downarrow$ SugSimGPT $\uparrow$
$MaxTokens=20$ 61.1% 37.92% 59.55% 75.54%
$MaxTokens=50$ 61.1% 44.39% 59.55% 71.40%
Transferability CodeCloak demonstrates full transferability between StartCoder and CodeLlama (i.e., training the DRL agent on StartCoder and using the trained agent on CodeLlama), which emphasizes its robustness and usefulness. The transferability results are presented in Table VI.
TABLE VI: Performance of CodeCloak in a cross-LLM transferability setup
Normalized Edit Distance ChatGPT Metrics
Prompts $\uparrow$ Suggestions $\downarrow$ LeakGPT $\downarrow$ SugSimGPT $\uparrow$
basemodel (StarCoder) 61.1% 37.92% 59.55% 75.54%
CodeLlma 13B 59.5% 35.42% 62.02% 74.53%
VI-F Evaluating CodeCloak on Complete Repositories
To assess the real-world effectiveness of CodeCloak we conducted experiments on complete repositories, simulating the entire development process utilizing a code assistant model, from start to finish. For each repository, we simulated (using our Developer Coding Simulator) the coding process twice: once without employing CodeCloak and again when utilizing our mitigation method.
When handling a sequence of prompts that arise during the process of developing a single codebase, the prompts sent to the code assistant exhibit a sequential dependency, where manipulations made to code segments in earlier prompts may impact subsequent prompts. To accurately replicate this sequential flow, we incorporated a preprocessing step before feeding the prompts to CodeCloak. This step involved checking the current manipulated prompt for any code segments that had previously been manipulated by CodeCloak within the same sequence (i.e., the same codebase) and applying the same manipulation to ensure consistency across the prompts.
After preprocessing, the prompt is fed to CodeCloak for further manipulation, aiming to minimize code leakage while preserving the relevance and usefulness of the code assistant’s suggestions. Finally, to quantify the degree of code leakage mitigation that CodeCloak achieves, we employed our code reconstruction method and reconstructed the codebase using the prompts captured during the simulated development process, both with and without CodeCloak.
To evaluate CodeCloak on complete repositories, we randomly select five repositories of different sizes from the 20 repositories evaluated in Sec. VI-D, and use the StarCoder model. CodeCloak demonstrates impressive performance: Without CodeCloak, the average code leakage rate across the evaluated repositories was 92.55%, indicating nearly complete code leakage. However, when employing CodeCloak, the leakage rate decreased dramatically to 44.3% on average, demonstrating a significant reduction in the code exposure. The SugSimGPT metric, which leverages ChatGPT’s code analysis capabilities, yielded an average similarity score of 71% between the original and manipulated suggestions, indicating a high level of preserved relevance. The edit distance metric, which measures visual similarity, obtained an average score of 39.2%, further demonstrating CodeCloak’s effectiveness in maintaining the usefulness of the code assistant’s suggestions throughout the development process.
TABLE VII: CodeCloak’s performance on complete repositories
Average Reconstruction Rate $\downarrow$ SugSimGPT $\uparrow$ Normalized Edit Distance Suggestions $\downarrow$
without CodeCloak 92.55% 100% 0%
with CodeCloak 44.30% 71.00% 39.20%
VII Conclusion
This work addressed the critical issue of code leakage risk arising from the increasing adoption of LLM-based code assistants. We proposed two complementary methods to mitigate this risk: (i) a code reconstruction technique for evaluating code exposure during development, and (ii) CodeCloak, a novel DRL agent that manipulates prompts sent to code assistant services. The leakage evaluation demonstrated significant code leakage risks with popular code assistants, underscoring the pressing need for effective mitigation strategies. CodeCloak addressed this risk successfully and mitigated the code leakage effectively while preserving the relevance and usefulness of the code assistant’s suggestions. These contributions pave the way for more secure and privacy-preserving utilization of LLM-based code assistants: Our code reconstruction method enables code leakage assessment, while CodeCloak offers a practical solution for mitigating leakage during the development process, facilitating responsible adoption of these powerful tools. We anticipate that our work will be the catalyst for further research addressing the security and privacy challenges associated with LLM-based code assistants.